Hoodmemes:

Chris buys 7 patatoes for 84 pesos, how many patatoes can he buy for 144 pesos?

Convert:

84/7= 12 144/12 = x 144/12 = 12 x=12 He can buy 12 potato's with 144 pesos.

Hoodmemes:

Ok,thank you.

Convert:

No problem! :)